Fermi National Laboratory
 

Accelerator Update for January 26, 2007 - January 29, 2007

 

Friday January 26

The day shift began with Operations monitoring the Tevatron (TeV) store 5197 (135.21E10) and the Recycler stash (4.5E10), with the Antiproton Source (Pbar) waiting for its vacuum to normalize before stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 8:14 AM, Operations reported that the Pelletron had tripped off. An operator and a Recycler expert investigated.

The Recycler expert had the Pelletron back on by 9:15 AM and then began doing electron cooling measurements.

Operations established beam to the Switchyard (SY) absorber at 9:22 AM.

Operations reported at 11:10 AM that vacuum experts were accessing the Pbar rings to pull out their vacuum pumps.

At 11:20 AM, Operations reported that the NuMI beam permit had tripped off due to a horn power supply airflow indication. Experts reset the supply, accessed the NuMI tunnel, inspected the air system and replaced air filters.

Operators search and secured the Meson MTest experimental areas at 11:44 AM.

Operations established beam to the Pbar Debuncher at 12:06 PM.

At 12:31 PM, a NuMI power supply (E:HV101) tripped off. It reset.

Operations reported at 1:08 PM that Pbar had begun stacking.

At 1:24 PM, a Booster RF station (BRF13) tripped off. Technicians bypassed the station and then began repairs.

At 1:58 PM, HV101 tripped off again.

At 2:02 PM, operators reported that they had reset HV101 a number of times. A power supply expert reported that he would replace the ground fault circuitry.

At 2:26 HV101 tripped again. An expert replaced the control board.

At 2:13 PM, a Linac RF station (LRF2) tripped off and wouldn’t reset. Linac experts investigated.

Operations reported that all of the Booster RF stations were up and running, except for BRF13. Experts were investigating.

Linac experts resolved the LRF2 problems. Operations resumed sending beam to Pbar and NuMI at 3:39 PM.

At 5:17 PM, a power supply expert reported that the HV101 ground fault problem was in the tunnel.

Operations prepared MI for an access at 5:28 PM.

At 6:30 PM, experts reported that they could see a LCW leak on the downstream end of HV101, but couldn’t tell where the leak was coming from due to the magnet’s shielding. This will be investigated further in the morning. NuMI was turned off. The first picture is of the leak and the second is the HV101 magnet and its shielding.

Booster RF experts reported at 7:34 PM that they had repaired a cable for BRF13.

Operations established beam to the Accumulator at 8:34 PM.

SY experts reported that they had completed the MTest beam commissioning for the evening.

At 11:27 PM, MIRF10 and BRF17 tripped off and wouldn’t reset. Operators investigated and then bypassed the stations. Operations called in experts.

Operations reported at midnight that the Recycler had only one stash of antiprotons for the evening shift.

Saturday January 27

The midnight shift began with Operations monitoring store 5197 (45.05E30) and the stash (34.5E10), and with Pbar stacking (52.4E10). NuMI was off due to a HV101 LCW leak.

A Booster RF expert reported that he replaced a 750V blocker circuit for BRF17. The station was back on by 12:52 AM. Operations established beam to the Pbar.

At 1:52 AM, the Linac Klystron Debuncher LCW pump tripped off and wouldn’t rest. Operations contacted an expert.

A MI RF expert reported at 1:53 AM that he had replaced some bad wiring for MIRF10.

Operations reported at 2:16 AM that they switched the Linac Debuncher LCW to its backup pump, per expert instruction.

Operations resumed sending beam to Pbar at 3:23 AM.

Operations reported at 8:13 AM that the Recycler had competed three stashes during the midnight shift.

At 8:39 AM, MIRF10 tripped off. Operators investigated and discovered that MIRF7 – 12 had all tripped off. The Duty Electrician reported that a yard breaker had tripped. He reset the breaker. Operations reported at 9:23 AM that MI was back on and that MIRF10 had been bypassed.

At 9:28 AM, Operations reported that the Pbar lithium lens power supply had tripped off due to an interlock trip: Collect Voltage Hi. After several resets, operators were able to turn the supply on at 9:37 AM and resumed stacking.

Operations terminated store 5197 at 10:18 AM.

Operations put the TeV into shot setup at 10:31 AM.

At 12:33 PM, after loading the protons and antiprotons from MI, Operations prepared MI and Recycler for an access.

Operations established store 5198 at 12:57 PM with an initial luminosity of 137.62E30.


Operations reported 3:53 PM that they found the NuMI LCW leak in MI-60. Experts began their repairs.

At 4:34 PM, vacuum technicians reported finding a MI vacuum leak on a dipole magnet between 851 and 852. They repaired the leak and the vacuum got better.

An expert reported at 6:54 PM that both NuMI LCW leaks were repaired. He estimated that it would take about two hours to replace the shielding for HV101.

A Booster expert completed an absorber study at 7:24 PM.

Operations reported at 8 PM that the Recycler had completed one stash since 8 AM.

At 8:58 PM, experts working on HV101 reported that the LCW leak had resumed and was worse that the initial leak. They valved out the magnet. The magnet must be replaced. The Run Coordinator said that this work would occur on Tuesday, 1/30/07.

Sunday January 28

The midnight shift began with Operations monitoring store 5198 and the stash, with Pbar stacking, and with NuMI off due to a HV101 LCW leak.

At 7:52 AM, Operations reported that a TeV RF station (TRF8) had tripped off.

Operations reported at 8 AM that the Recycler had stashed four times in the last 12 hours.

A SY expert began a study at 9:35 AM.

At 12:39 PM, a MI expert reported that coalescing was bad. He found that MIRF13 was bad. Operators bypassed the station.

At 12:51 PM, Operations noticed that a Pbar control program had set a Debuncher power supply (D:VT714) to its limit, but its readback showed zero amps. Operators investigated and the supply began working normally all on its own.

Operations terminated store 5198 at 5:01 PM.

Operations put the TeV into shot setup at 5:14 PM.

Operations established store 5199 at 7:12 PM with an initial luminosity of 196.31E30.

Operations reported at 8 PM that the Recycler had stashed four times over the last 12 hours.

At 11:25 PM, an Accumulator RF station (ARF1) tripped off. Operators investigated and then reset the station.

Monday January 29

The midnight shift began with Operations monitoring store 5199 and the stash, with Pbar stacking, and with NuMI off.

At 5:06 AM, LRF2 tripped off. It reset okay.

Operations reported at 8 AM that the Recycler had stashed five times during the last 12 hours.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV, MI, and Pbar experts to conduct studies. The HV101 magnet replacement will began after the next store is established, which should be around 4 AM on Tuesday, 1/30/07. The repair will take about 72 hours. MI, Pbar, SY, and NuMI will be off until Friday, 2/2/07.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/29/2007    email Fermilab